OpenStudy (anonymous):

how will this be solved 4/3(12xy) ??/ pls help me out

OpenStudy (anonymous):

how can this be answered??

OpenStudy (anonymous):

Solving for x, we have: 4/(3(12xy)) = 0? But that does not help at all. Need more information pleaes.

OpenStudy (anonymous):

how will i multiply it...??

OpenStudy (anonymous):

What is the context?

OpenStudy (anonymous):

what du u meen?

OpenStudy (anonymous):

Like I said above this equation is equal to saying 4/36xy = 0. If multiply across we have: 4 = 0 as anything times 0 is zero. What is the subject you are studying and what is the lesson called?

jhonyy9 (jhonyy9):

4/3(4*3xy)=16xy - if 12xy is in numerator and simplify by3